OpenStudy (anonymous):

If the sin 60° is square root 3 over 2, then the cos ____° = _____. 30° ; 1 over 2 60° ; square root 2 over 2 30° ; square root 3 over 2 60° ; 1

OpenStudy (anonymous):

sin(\theta) = cos(90-\theta). Need more help?

OpenStudy (anonymous):

 C